# Matteo Escudé

### v5

@@ 4.3 More verifiability may hurt the sender
At first glance, moving to a $\succeq^{\mathrm{lc}}$-higher (or, perhaps, even $\succeq$-higher, as introduced in Remark 3) structure might appear (via a standard replication argument) to be always beneficial for the sender, regardless of whether she can prove news better than the prior. This is not the case, as highlighted by the following example.
Fix some finite set of messages $\tilde{M}$ and some message $m_{H}\notin\tilde{M}$. Let verifiability structure $M^{\prime}$ be such that $M^{\prime}(s)=\tilde{M}$ for each $s\in[0,1]$, while $M^{\prime\prime}$ is such that $M^{\prime\prime}(s)=\tilde{M}$ for $s<9/10$ and $M^{\prime\prime}(s)=\tilde{M}\cup\{m_{H}\}$ for $s\geq 9/10$. $M^{\prime}$ is therefore a cheap-talk structure (as introduced in Example 3). $M^{\prime\prime}$ lets the sender to use any of the cheap-talk messages in $\tilde{M}$, but also allows her to prove particularly good news (namely: if she discovers that the state is more than 90% likely to be $1$, she can prove it by sending message $m_{H}$). It is immediate that $M^{\prime\prime}\succeq M^{\prime}$ (and thus $M^{\prime\prime}\succeq^{\mathrm{lc}}M^{\prime}$). [^29] This is an example of an ‘adding a message’ verifiability shift from Example 7. Let $v$ be defined as follows:
$v(s)=\begin{cases}0\quad\text{if}\quad s<2/5,\\ 2\quad\text{if}\quad s\in[2/5,4/5),\\ 3\quad\text{if}\quad s\geq 4/5.\end{cases}$
I now argue, applying directly the results from the previous section, that the (unique) equilibrium payoff with structure $M^{\prime\prime}$ is strictly lower than the (unique) equilibrium payoff with structure $M^{\prime}$. Observe first that, since under $M^{\prime}$ the sender cannot prove news better than the prior, Proposition 1 implies that the sender’s highest expected payoff in equilibrium is $v(\overline{p})=2$ (in this example, it is immediate that this is the sender’s unique equilibrium payoff). Under $M^{\prime\prime}$ however, the sender can prove news better than the prior, so Corollary 2 implies that the sender’s equilibrium payoff is $5/3<2$. This is illustrated in Figure 2.
@@
Figure 2: The bold line represents the sender’s payoff function $v$. In each case, the sender’s expected payoff in equilibrium is represented by the large gray dot. In the right-hand panel (verifiability structure $M^{\prime\prime}$) the small gray dots represent the two ex post sender payoffs that occur with positive probability in equilibrium.
The reason why a replication argument fails here is that the sender cannot commit to acquiring no information under $M^{\prime\prime}$, thereby replicating her payoff under $M^{\prime}$. Informally, this is the case because ‘claiming ignorance’ is credible under $M^{\prime}$ (using any message in $\tilde{M}$), but it is not under $M^{\prime\prime}$. That’s because under $M^{\prime\prime}$ the sender can prove good news (with $m_{H}$), so she will seek them, only claiming that ‘she doesn’t know’ (using a message in $\tilde{M}$) when she did not find said good news. Increasing separation possibilities may therefore worsen the meaning of messages in equilibrium (in this case: any message in $\tilde{M}$ must lead to a payoff of 00 under $M^{\prime\prime}$, while it lead to a payoff of $2$ under $M^{\prime}$). This can, in turn, lower the equilibrium payoff for the sender.
@@ 4.4 Value of commitment and sender-optimal verifiability
We showed that increasing verifiability, in the sense of enlarging the set of lowest-consistent signal realizations, is both necessary and sufficient to make the sender ex ante better off, provided that we are considering environments in which she can prove news better than the prior.
In light of this result, it is straightforward to characterize the structures that are optimal for the sender, in the sense that the equilibrium payoff is highest among all admissible ones. Observe first that Corollary 2 implies that the sender can obtain her full-commitment payoff $(\operatorname*{cav}v)(\overline{p})$ (that is: the payoff she would obtain under public information acquisition, as in Kamenica and Gentzkow (2011)) with the ‘mandatory disclosure’ verifiability structure (the one of Example 1: $M(s)=\{s\}$, $s\in[0,1]$). Therefore, characterizing sender-optimal structures is tantamount to characterizing structures that allow her to obtain such full-commitment payoff:
@@
  1. (1)
    In every equilibrium the sender obtains the full-commitment payoff for any admissible $v$ and $\overline{p}$.
  2. (2)
    All $s\in[0,1]$ are lowest-consistent.
The characterization therefore highlights precisely in which sense the full commitment assumption in Kamenica and Gentzkow (2011) can be relaxed without affecting the sender’s equilibrium payoff. Provided that for every sender type there is a feasible message that no lower type can send, her equilibrium expected payoff matches the full-commitment one. If instead some interim type is not lowest-consistent, it is possible to find a set of admissible parameters such that her equilibrium expected payoff is strictly lower than the full-commitment one.
@@
Let $M^{\textit{md}}$ denote the mandatory disclosure verifiability structure ($M^{\textit{md}}(s)=\{s\}$, $s\in[0,1]$). I first show that (1) implies (2). Suppose that some $s\in[0,1]$ is not lowest-consistent with structure $M$. Then $M\not\succeq^{\mathrm{lc}}M^{\textit{md}}$. It follows from Theorem 2(b) that there exist admissible $v$ and $\overline{p}$ such that the sender obtains a payoff strictly below the full-commitment one, which is the unique outcome with $M^{\textit{md}}$.
I show next that (2) implies (1). If $v(\overline{p})=v(1)$ the result is immediate, since the sender obtains the maximal payoff of $v(\overline{p})$ in every equilibrium. If $v(\overline{p})<v(1)$, the sender can prove news better than the prior with $M^{\textit{md}}$. Consider any structure $M$ such that all $s\in[0,1]$ are lowest-consistent. Then $M\succeq^{\mathrm{lc}}M^{\textit{md}}$, so she must obtain at least the same payoff with $M$ as she does with $M^{\textit{md}}$ by Theorem 2(a). This payoff cannot be strictly higher than the full-commitment one, so they must coincide.
@@ 5 Receiver-optimal verifiability
Which verifiability structures are desirable for the receiver? I show that a simple verifiability structure with two messages is such that the receiver obtains full information in equilibrium. Provided that she values information about the state, this structure will therefore be optimal for her across all admissible structures. [^30] In more detail: if the receiver obtains a payoff $u(p)$ at belief $p$, obtaining full information about the state is optimal (across all information structures) if $u$ is convex. That is the case, for example, if $u(p)$ is the value of a decision problem under uncertainty with expected-utility preferences i.e. if there is an action set $A$ and a continuous function $f:A\times\{0,1\}\to\mathbb{R}$ such that $u(p)=\max_{a\in A}pf(a,1)+(1-p)f(a,0)$ for each $p\in[0,1]$. I also provide a converse: only structures in a specific binary class are such that, no matter what the (admissible) preferences and prior of the sender, she will provide full information in every equilibrium.
Proposition 3 (Receiver-optimal verifiability).
@@
  1. (1)
    In every equilibrium the sender provides full information for any $(v,\overline{p})$ such that $v(\overline{p})<v(1)$.
  2. (2)
    Only types 00 and $1$ are lowest-consistent.
$v(\overline{p})<v(1)$ is a very mild requirement on what value the sender can potentially obtain from persuasion. If otherwise $v(\overline{p})=v(1)$—recall that $v$ is non-decreasing—the sender would be obtaining her highest possible payoff at the prior so, unsurprisingly, equilibria in which the sender does not provide full information persist. [^31] If $v(\overline{p})=v(1)$ and only types 00 and $1$ are lowest-consistent, it can be shown that there exists an equilibrium in which the sender provides full information. However, for the sender such an equilibrium may be (substantially) worse than the no-information equilibrium, which always exists and is sender-optimal, as she cannot prove news better than the prior (this follows from Proposition 1).
The class receiver-optimal structures takes an intuitive form. It can be interpreted as the receiver committing to only checking the sender’s claim that ‘news are good’ if the sender produces evidence that news are conclusively good (i.e. if the sender proves that the state is $1$). In this situation, the only hope the sender has of swaying the receiver in her favor is to learn with maximal probability that the state is $1$, which is achieved by also learning that the state is 00 with certainty. As all acquired information is transmitted to the receiver (from Corollary 1), the latter also learns the value of the state. For the receiver, committing to checking any other intermediate claim is potentially sub-optimal, as it provides the sender more commitment power.
@@
Consider first (2) $\Rightarrow$ (1). If $1$ is lowest-consistent and $v(\overline{p})<v(1)$ then the sender can prove news better than the prior. It follows from Theorem 1 that the only candidate equilibrium involves the sender acquiring full information (since only 00 and $1$ are lowest-consistent). In such an equilibrium, the receiver perfectly learns the value of the state (from Corollary 1). Since an equilibrium exists (Appendix A), the sender provides full information in equilibrium.
Consider now (1) $\Rightarrow$ (2). Suppose that some $s^{*}\notin\{0,1\}$ is also lowest-consistent and let $v(s)=1(s\geq s^{*})$. Let $\overline{p}\equiv s^{*}/2$. Corollary 2 implies that the sender’s unique equilibrium payoff with is $\overline{p}/s$, which is obtained by splitting the prior on $\{0,s\}$. Hence the sender acquires and provides less-than-full information.
@@ A.1 Equilibrium definition
    for all $\pi^{\prime}\in\Pi$.
  2. (2)
    Sequentially rational communication. The sender chooses an optimal message: $v(\beta(m))\geq v(\beta(m^{\prime}))$ for every $m\in\operatorname*{supp}\mu(\cdot|\pi^{\prime},s)$ and any $m^{\prime}\in M(s)$, for each $\pi^{\prime}\in\Pi$ and $s\in[0,1]$.
  3. (3)
    Consistent receiver beliefs. Beliefs are consistent with the verifiability structure: $\beta(m)\in\operatorname*{conv}M^{-1}(m)$ for all $m\in\mathcal{M}$. If $m\in\mathcal{M}$ is on-path, it is further required that
@@ A.2.1 Preliminaries
If $w_{\beta}$ is upper semi-continuous, a sender-optimal signal exists and attains value $(\operatorname*{cav}w_{\beta})(\overline{p})$. If signal $\pi$ is such that $\operatorname*{supp}\pi_{S}=\{s_{-}(\overline{p},w_{\beta}),s_{+}(\overline{p},w_{\beta})\}$ then it is sender-optimal.
@@
$\operatorname*{cav}w_{\beta}$ is affine over $[s_{-}(\overline{p},w_{\beta}),s_{+}(\overline{p},w_{\beta})]$.
In what follows we will frequently use the always skeptical belief function, defined as $\underline{\beta}(m)\equiv\min M^{-1}(m)$ for each message $m\in\mathcal{M}$ [^32] The right-hand side is well-defined as $M^{-1}(\cdot)$ is assumed to be closed (see the technical assumptions in Section 2.1.1).
A.2.2 Existence proof
The existence proof is divided into two cases: Lemma 3 and Lemma 4, dealing separately with the cases in which the sender can or cannot prove news better than the prior (Definition 3). Lemma 3 shows more than mere existence as it constructs a specific equilibrium in which the sender acquires no information, which is used to prove part of Proposition 1.
@@ A.2.2 Existence proof
If the sender cannot prove news better than the prior, an equilibrium in which the sender acquires no information exists.
@@
  1. (1)
    $\pi\in\Pi$ is such that $\pi_{S}(\overline{p})=1$.
  2. (2)
    $\mu(m|\pi^{\prime},s)=1$ for any one arbitrary $m\in\operatorname*{arg\,max}_{m\in M(s)}\min M^{-1}(m)$ for every $\pi^{\prime}\in\Pi$ and each $s\in[0,1]$ for which $m_{0}\notin M(s)$.
    $\mu(m_{0}|\pi^{\prime},s)=1$ for any $\pi^{\prime}\in\Pi$ and $s\in[0,1]$ for which $m_{0}\in M(s)$.
  3. (3)
    $\beta(m)=\underline{\beta}(m)$ except for some arbitrary $m_{0}\in M(\overline{p})$ for which we impose that $\beta(m_{0})=\overline{p}$.
Observe first that, by construction, the sender’s messaging strategy is optimal whenever $m_{0}$ is not available and that the receiver’s beliefs are consistent. It remains to show that for the sender it is optimal to acquire no information (point (1) above) and send message $m_{0}$ at signal realizations where $m_{0}$ is available. Observe that, since the sender cannot prove news better than the prior, $\beta(m)\leq\beta(m_{0})$ for all $m\in\mathcal{M}$. Since $v$ is non-decreasing, it follows that $v(\beta(m))\leq v(\beta(m_{0}))$ for all $m\in\mathcal{M}$. This in turn implies both that sending $m_{0}$ is optimal, when feasible, and that acquiring no information, as specified by (1), is an optimal information acquisition strategy, as message $m_{0}$ is then sent with probability 1.
Lemma 4 (Existence with PNBP).
If the sender can prove news better than the prior, an equilibrium exists.
@@
The candidate equilibrium $(\pi,\mu,\beta)$ is defined as follows.
  1. (1)
    $\pi\in\Pi$ is such that $\operatorname*{supp}\pi_{S}=\{s_{-}(\overline{p},w_{\beta}),s_{+}(\overline{p},w_{\beta})\}$, where $w_{\beta}$, $s_{-}$ and $s_{+}$ are defined in equations (3), (4) and (5).
  2. (2)
    $\mu(m|\pi^{\prime},s)=1$ for any one arbitrary $m\in\operatorname*{arg\,max}_{m\in M(s)}\min M^{-1}(m)$ for every $\pi^{\prime}\in\Pi$ and each $s\in[0,1]$.
  3. (3)
    $\beta(m)=\underline{\beta}(m)$ for all $m\in\mathcal{M}$.
To streamline notation, write $s_{-}$ ($s_{+}$) to denote $s_{-}(\overline{p},w_{\beta})$ ($s_{+}(\overline{p},w_{\beta})$) as we will be holding $\overline{p}$, $\beta$ and $w_{\beta}$ fixed throughout the rest of the proof. Suppose first that both $s_{-}$ and $s_{+}$ are lowest-consistent (Definition 2). Then, from the definition of $\mu$, $s_{-}$ ($s_{+}$) is lowest-consistent with any $m\in\operatorname*{supp}\mu(m|\pi,s_{-})$ ($m\in\operatorname*{supp}\mu(m|\pi,s_{+})$). By construction, the receiver’s belief is therefore correct following on-path messages and consistent off-path. Given $\beta$, the sender’s messaging strategy is optimal by construction. Observe finally that the technical assumptions at the end of §2.1 ensure that $w_{\beta}$ is upper semi-continuous, hence Lemma 1 implies that $\pi$ is an optimal signal. It follows that, provided that both $s_{-}$ and $s_{+}$ are lowest-consistent, the candidate is indeed an equilibrium.
To complete the proof we must therefore show that $s_{-}$ and $s_{+}$ are lowest-consistent. The argument is divided into two separate cases .
@@
If $s_{-}=s_{+}=\overline{p}$, then $s_{-}$ and $s_{+}$ are lowest-consistent.

### v6

@@ 4.3 More verifiability may hurt the sender
At first glance, moving to a $\succeq^{\mathrm{lc}}$-higher (or, perhaps, even $\succeq$-higher, as introduced in Remark 3) structure might appear (via a standard replication argument) to be always beneficial for the sender, regardless of whether she can prove news better than the prior or whether an unraveling equilibrium exists. The following example demonstrates that this is not the case: the mere existence of a verifiable ‘good news’ message can taint the meaning of ‘claiming ignorance’, turning the inability to produce good news into bad news—thereby hurting the sender.
Fix some finite set of messages $\bar{M}$ and some message $m_{E}\notin\bar{M}$. Let verifiability structure $M^{\prime}$ be such that $M^{\prime}(s)=\bar{M}$ for each $s\in[0,1]$, while $M^{\prime\prime}$ is such that $M^{\prime\prime}(s)=\bar{M}$ for $s<0.9$ and $M^{\prime\prime}(s)=\bar{M}\cup\{m_{E}\}$ for $s\geq 0.9$. $M^{\prime}$ is therefore a cheap talk structure (as introduced in Example 3). $M^{\prime\prime}$ lets the sender use any of the cheap talk messages in $\bar{M}$, but also allows her to prove excellent news (namely: if she discovers that the state is more than 90% likely to be $1$, she can prove it by sending message $m_{E}$). It is immediate that $M^{\prime\prime}\succeq M^{\prime}$ (and thus $M^{\prime\prime}\succeq^{\mathrm{lc}}M^{\prime}$).
Finally, let $v$ be defined as follows: $v(p)=0$ for $p<0.4$, $v(p)=2$ for $p\in[0.4,0.8)$ and $v(p)=3$ for $p\geq 0.8$. Fix the prior to be $\overline{p}=1/2$.
I now argue, applying directly the results from Section 3, that the (unique) equilibrium payoff with structure $M^{\prime\prime}$ is strictly lower than the (unique) equilibrium payoff with structure $M^{\prime}$.
Since under $M^{\prime}$ the sender cannot prove news better than the prior, Theorem 2 implies that the sender’s highest expected payoff in equilibrium is $v(\overline{p})=2$ (in this example, it is immediate that this is the sender’s unique equilibrium payoff), as illustrated in Figure 3(a) for the case in which there is a single cheap talk message: $\bar{M}=\{m_{\emptyset}\}$. Under $M^{\prime\prime}$ however, the sender can prove news better than the prior, so Theorem 2 and Corollary 1 together imply that the sender’s equilibrium expected payoff is $5/3<2$, as illustrated in Figure 3(b).
@@
Figure 3: On the left: the sender value function (solid black)—overlaid over $v$ (solid gray). On the right: the sender’s skepticism-adjusted value function (solid black)—overlaid over $v$ (solid gray)—and its concave envelope (thick, dotted). In each case, the sender’s equilibrium expected payoff is represented by the large black dot.
The reason why a replication argument fails here is that the sender cannot commit to acquiring no information under $M^{\prime\prime}$, thereby replicating her payoff under $M^{\prime}$. Informally, this is the case because ‘claiming ignorance’ is credible under $M^{\prime}$ (using any cheap talk message in $\bar{M}$), but it is not under $M^{\prime\prime}$. This is because under $M^{\prime\prime}$ the sender can prove good news (with $m_{E}$), so she will seek it, only claiming ignorance (using a message in $\bar{M}$) when she did not find said good news. Increasing separation possibilities may therefore worsen the meaning of messages in equilibrium (in this case: any message in $\bar{M}$ must lead to a payoff of 0 under $M^{\prime\prime}$, while it led to a payoff of $2$ under $M^{\prime}$). This can, in turn, lower the equilibrium expected payoff for the sender.
@@ 4.4 Value of commitment and sender-optimal verifiability
We showed that increasing verifiability, in the sense of enlarging the set of sender types that are lowest-consistent with some message, is both necessary and sufficient to make the sender ex ante weakly better off, provided that we are considering environments in which she can prove news better than the prior.
In light of this result, it is straightforward to characterize the structures that are optimal for the sender, in the sense that her equilibrium expected payoff is highest among all admissible ones. Note first that the sender can obtain her full-commitment payoff $(\operatorname*{cav}v)(\overline{p})$ (that is: the payoff she would obtain under public information acquisition, as in Kamenica and Gentzkow (2011)) with the ‘mandatory disclosure’ verifiability structure (the one of Example 1: $M(s)=\{s\}$, $s\in[0,1]$). This follows from Theorem 2 and Corollary 1, if she can prove news better than the prior, and directly from Theorem 2, if she cannot. [^23] In more detail: if with the mandatory disclosure verifiability structure the sender cannot prove news better than the prior, it must mean that $v(\overline{p})=\max_{p\in[0,1]}v(p)$, so that $(\operatorname*{cav}v)(\overline{p})=v(\overline{p})$. This payoff is therefore attained in an unpersuasive equilibrium, which exists by Theorem 2.
Therefore, characterizing sender-optimal structures is tantamount to characterizing structures that allow her to obtain such full-commitment payoff:
@@
  1. (i)
    In every equilibrium the sender obtains the full-commitment payoff for every admissible $v$ and $\overline{p}$.
  2. (ii)
    Every $s\in[0,1]$ is lowest-consistent with some message.
The characterization therefore highlights precisely in which sense the full commitment assumption in Kamenica and Gentzkow (2011) can be relaxed without affecting the sender’s equilibrium payoff. Provided that for every sender type there is a feasible message that no lower type can send, her equilibrium expected payoff matches the full-commitment one. If instead some type is not lowest-consistent with any message, it is possible to find a set of admissible parameters such that her equilibrium expected payoff is strictly lower than the full-commitment one.
@@
Let $M^{\textit{md}}$ denote the mandatory disclosure verifiability structure ($M^{\textit{md}}(s)=\{s\}$, $s\in[0,1]$). Consider first (i) $\Rightarrow$ (ii). Suppose that $M$ is such that some $s\in[0,1]$ is not lowest-consistent with any message. Then $M\not\succeq^{\mathrm{lc}}M^{\textit{md}}$. It follows from Theorem 3 that there exist admissible $v$ and $\overline{p}$ such that the sender obtains a payoff strictly below the full-commitment one, which is the unique outcome with $M^{\textit{md}}$.
Consider now (ii) $\Rightarrow$ (i). If $v(\overline{p})=v(1)$ the result is immediate, since the sender obtains the maximal payoff of $v(\overline{p})$ in every equilibrium. If $v(\overline{p})<v(1)$, the sender can prove news better than the prior with $M^{\textit{md}}$. Consider any structure $M$ such that each $s\in[0,1]$ is lowest-consistent with some message. The sender can also prove news better than the prior with $M$, so her expected equilibrium payoff is unique (this follows from Theorem 2 and Corollary 1). Since $M\succeq^{\mathrm{lc}}M^{\textit{md}}$, she must obtain at least the same payoff with $M$ as she does with $M^{\textit{md}}$ by Theorem 3. This payoff cannot be strictly higher than the full-commitment one, so they must coincide. ∎
@@ 5 Receiver-optimal verifiability
Which verifiability structures are desirable for the receiver? I show that a simple verifiability structure with two messages is such that the sender provides full information in equilibrium. [^24] I say that the sender provides full information in equilibrium if, conditional on $\omega=1$, the receiver holds a posterior belief of $1$ (with probability one) and, conditional on $\omega=0$, the receiver holds a posterior belief of 0 (with probability one). Provided that the receiver values information about the state, this structure will therefore be optimal for her across all admissible structures. [^25] In more detail: if the receiver obtains a payoff $u(p)$ at belief $p$, obtaining full information about the state is optimal (across all information structures) if $u$ is convex. That is the case, for example, if $u(p)$ is the value of a decision problem under uncertainty with expected-utility preferences: there is an action set $A$ and a continuous function $f:A\times\{0,1\}\to\mathbb{R}$ such that $u(p)=\max_{a\in A}pf(a,1)+(1-p)f(a,0)$ for each $p\in[0,1]$. I also provide a converse: only structures in a specific binary class are such that, no matter what the (admissible) preferences and prior of the sender, she will provide full information in every equilibrium.
Proposition 2 (Full information provision).
@@
  1. (i)
    In every equilibrium the sender provides full information for every $(v,\overline{p})$ such that $v(\overline{p})<v(1)$.
  2. (ii)
    Only types 0 and $1$ are lowest-consistent with some message.
$v(\overline{p})<v(1)$ is a very mild requirement on what value the sender can potentially obtain from persuasion. If otherwise $v(\overline{p})=v(1)$—recall that $v$ is non-decreasing—the sender would be obtaining her highest possible payoff at the prior so, unsurprisingly, equilibria in which the sender does not provide full information persist. [^26] If $v(\overline{p})=v(1)$ and only types 0 and $1$ are lowest-consistent with some message, it can be shown that there exists an equilibrium in which the sender provides full information. However, for the sender such an equilibrium may be (substantially) worse than an unpersuasive equilibrium, which always exists and is sender-optimal, as she cannot prove news better than the prior (this follows from Theorem 2).
The class of verifiability structures leading to full information provision takes an intuitive form. It can be interpreted as the receiver committing to only checking the sender’s claim of ‘good news’ if the latter produces evidence that news is conclusively good (i.e., if the sender proves that the state is $1$).
In this situation, the only hope the sender has of swaying the receiver in her favor is to learn with maximal probability that the state is $1$, which is achieved by also learning that the state is 0 with certainty. Since type $1$ is lowest-consistent with some message and $v(\overline{p})<v(1)$, the sender can prove news better than the prior. Therefore, Theorem 2 implies that all acquired information is revealed to the receiver, so the latter also learns the value of the state. For the receiver, committing to checking any other intermediate claim provides the sender with more commitment power, so may lead to less-than-full information provision.
@@
Consider first (ii) $\Rightarrow$ (i). Observe that the sufficient condition stated in Lemma 1 is satisfied, so an equilibrium exists. If type $1$ is lowest-consistent with some message and $v(\overline{p})<v(1)$ then the sender can prove news better than the prior, so all equilibria must be unraveling (Theorem 2). Therefore, in equilibrium, $\pi_{S}(1)=\overline{p}$ and $\pi_{S}(0)=1-\overline{p}$, since only 0 and $1$ are lowest-consistent with some message. In any unraveling equilibrium the sender reveals all acquired information (Theorem 1), hence the sender provides full information.
Consider now (i) $\Rightarrow$ (ii). Suppose that some $s^{*}\notin\{0,1\}$ is also lowest-consistent with some message. Let $v(p)=1_{\{p\geq s^{*}\}}$ and $\overline{p}=s^{*}/2$. The following is an unraveling equilibrium in which the sender acquires and provides less-than-full information.
The receiver is maximally skeptical, assigning belief $\beta(m)\equiv\min M^{-1}(m)$ to every $m\in\mathcal{M}$. The sender chooses a signal that splits the prior on $\{0,s^{*}\}$, sending any available message when 0 realizes and sending a message that type $s^{*}$ is lowest-consistent with when $s^{*}$ realizes. Letting $w_{\beta}(s)\equiv\max_{m\in M(s)}v(\beta(m))$, $s\in[0,1]$, observe that the sender’s strategy attains an expected payoff equal to $(\operatorname*{cav}w_{\beta})(\overline{p})=(\operatorname*{cav}v)(\overline{p})=1/2$, and is therefore optimal. The receiver’s beliefs are updated using Bayes’ rule on-path, and consistent with the verifiability structure by construction. This profile is therefore an unraveling equilibrium, as required. ∎
@@ A.1 Equilibrium definition
    for every $\pi^{\prime}\in\Pi$.
  2. (ii)
    Sequentially rational communication. The sender chooses an optimal message: $v(\beta(m))\geq v(\beta(m^{\prime}))$ for every $m\in\operatorname*{supp}\mu(\cdot|\pi^{\prime},s)$ and every $m^{\prime}\in M(s)$, for every $\pi^{\prime}\in\Pi$ and $s\in[0,1]$.
  3. (iii)
    Consistent receiver beliefs. Beliefs are consistent with the verifiability structure: $\beta(m)\in\operatorname*{conv}M^{-1}(m)$ for every $m\in\mathcal{M}$. If $m\in\mathcal{M}$ is on-path, it is further required that
@@ A.2.1 Preliminaries
If $w_{\beta}$ is upper semi-continuous, a sender-optimal signal exists and attains value $(\operatorname*{cav}w_{\beta})(\overline{p})$. The signal $\pi$ with $\operatorname*{supp}\pi_{S}=\{s_{-}(\overline{p},w_{\beta}),s_{+}(\overline{p},w_{\beta})\}$ is sender-optimal.
@@
$\operatorname*{cav}w_{\beta}$ is affine over $s_{-}(\overline{p},w_{\beta}),s_{+}(\overline{p},w_{\beta})]$.
A.2.2 Proof of Lemma 1
The existence proof is divided into two parts: Lemma 5 and Lemma 6, dealing separately with the cases in which the sender can or cannot prove news better than the prior (Definition 5). Lemma 5 shows more than mere existence as it constructs an unpersuasive equilibrium in which the sender acquires no information, which is used to prove the second part of Theorem 2. Lemma 5 does not rely on the continuity assumption on $M$, namely upper semi-continuity of $s\mapsto\max_{m\in M(s)}\min M^{-1}(m)$, for $s\in[0,1]$, stated in Lemma 1. Lemma 6 does rely on it.
@@ A.2.2 Proof of Lemma 1
If the sender cannot prove news better than the prior, an unpersuasive equilibrium in which the sender acquires no information exists.
Proof.
Fix some $m_{0}\in M(\overline{p})$. The candidate equilibrium $(\pi,\mu,\beta)$ is defined as follows:
  1. (i)
    The signal $\pi$ is such that $\pi_{S}(\overline{p})=1$.
  2. (ii)
    $\mu(m|\pi^{\prime},s)=1$ for some (arbitrary) $m$ in $\operatorname*{arg\,max}_{m\in M(s)}\min M^{-1}(m)$, for every $\pi^{\prime}\in\Pi$ and $s\in[0,1]$ for which $m_{0}\notin M(s)$; $\mu(m_{0}|\pi^{\prime},s)=1$ for every $\pi^{\prime}\in\Pi$ and $s\in[0,1]$ for which $m_{0}\in M(s)$.
  3. (iii)
    $\beta(m)=\min M^{-1}(m)$ if $m\neq m_{0}$ and $\beta(m_{0})=\overline{p}$.
Observe first that, by construction, the sender’s messaging strategy is optimal whenever $m_{0}$ is not available (since $v$ is non-decreasing) and that the receiver’s beliefs are consistent. It remains to show that for the sender it is optimal to acquire no information (item (i) above) and send message $m_{0}$ at signal realizations where $m_{0}$ is available. Observe that, since the sender cannot prove news better than the prior and $v$ is non-decreasing, $v(\beta(m))\leq v(\beta(m_{0}))$ for all $m\in\mathcal{M}$. This in turn implies both that sending $m_{0}$ is optimal, when feasible, and that acquiring no information, as specified by item (i), is an optimal information acquisition strategy, as message $m_{0}$ is then sent with probability one. ∎
Lemma 6 (Existence with PNBP).
If the sender can prove news better than the prior and $s\mapsto\max_{m\in M(s)}\min M^{-1}(m)$, $s\in[0,1]$ is upper semi-continuous, an equilibrium exists.
@@
  1. (i)
    The signal $\pi$ is such that $\operatorname*{supp}\pi_{S}=\{s_{-}(\overline{p},w_{\beta}),s_{+}(\overline{p},w_{\beta})\}$, where $w_{\beta}$, $s_{-}$ and $s_{+}$ are defined in equations (4), (5) and (6).
  2. (ii)
    $\mu(m|\pi^{\prime},s)=1$ for some (arbitrary) $m$ in $\operatorname*{arg\,max}_{m\in M(s)}\min M^{-1}(m)$ for every $\pi^{\prime}\in\Pi$ and $s\in[0,1]$.
  3. (iii)
    $\beta(m)=\min M^{-1}(m)$ for every $m\in\mathcal{M}$.
To streamline notation, write $s_{-}$ ($s_{+}$) to denote $s_{-}(\overline{p},w_{\beta})$ ($s_{+}(\overline{p},w_{\beta})$) as we will be holding $\overline{p}$, $\beta$ and $w_{\beta}$ fixed throughout the rest of the proof. Suppose first that $s_{-}$ ($s_{+}$) is lowest-consistent (Definition 2) with some message. Then, it follows from the definition of $\mu$, that $s_{-}$ ($s_{+}$) is lowest-consistent with every $m\in\operatorname*{supp}\mu(m|\pi,s_{-})$ ($m\in\operatorname*{supp}\mu(m|\pi,s_{+})$).
By construction, the receiver’s belief is therefore correct following on-path messages and consistent off-path. Given $\beta$, the sender’s messaging strategy is optimal by construction, as it maximizes the receiver’s posterior belief and $v$ is non-decreasing. Finally, notice that $w_{\beta}$ is upper semi-continuous. This follows from two observations. First,
$w_{\beta}(s)=\max_{m\in M(s)}v\big(\min M^{-1}(m)\big)=v\bigg(\max_{m\in M(s)}\min M^{-1}(m)\bigg),$
where the first equality holds by construction (since $\beta=\min M^{-1}$) and the second equality follows from $v$ being non-decreasing. Second, $s\mapsto v\big(\max_{m\in M(s)}\min M^{-1}(m)\big)$ is upper semi-continuous, since $v$ is upper semi-continuous and non-decreasing and $s\mapsto\max_{m\in M(s)}\min M^{-1}(m)$ is upper semi-continuous by hypothesis.
Therefore, since $w_{\beta}$ is upper semi-continuous, Lemma 3 implies that $\pi$ is an optimal signal. It follows that, provided that each of $s_{-}$ and $s_{+}$ is lowest-consistent with some message, the candidate is indeed an equilibrium.
To complete the proof we must therefore show that $s_{-}$ and $s_{+}$ are indeed lowest-consistent with some message. The argument is divided into two separate cases.
Claim 1.
If $s_{-}=s_{+}=\overline{p}$, then $s_{-}$ and $s_{+}$ are lowest-consistent with some message.
@@
If $s_{-}=s_{+}=\overline{p}$ then it must be that $(\operatorname*{cav}w_{\beta})(\overline{p})=w_{\beta}(\overline{p})$, by definition of $s_{-}$ and $s_{+}$. Suppose otherwise that $\overline{p}$ is not lowest-consistent with any message. Then there exists some $\tilde{s}<\overline{p}$ such that $w_{\beta}(\tilde{s})\geq w_{\beta}(\overline{p})$, since any message available at $\overline{p}$ (including optimal ones) is also available at some lower type. Recall also that, since the sender can prove news better than the prior, there exists some message $m_{h}$ such that $w_{\beta}(\min M^{-1}(m_{h}))>w_{\beta}(\overline{p})$ with $s_{h}\equiv\min M^{-1}(m_{h})>\overline{p}$, since $v$ is non-decreasing. Summarizing, we have that $\tilde{s}<\overline{p}<s_{h}$ with $w_{\beta}(\tilde{s})\geq w_{\beta}(\overline{p})$ and $w_{\beta}(s_{h})>w_{\beta}(\overline{p})$. Since $\operatorname*{cav}w_{\beta}$ is concave and majorizes $w_{\beta}$, it must be that $(\operatorname*{cav}w_{\beta})(\overline{p})>w_{\beta}(\overline{p})$, contradicting that $(\operatorname*{cav}w_{\beta})(\overline{p})=w_{\beta}(\overline{p})$. ∎
@@
If $s_{-}<\overline{p}<s_{+}$, then $s_{-}$ and $s_{+}$ are lowest-consistent with some message.